OpenStudy (anonymous):

So here is the informatiovn: A delivery serice charges a base price for an overnight delivery of a package plus an extra charge for each pound the package weighs. A customer is billed $22.85 for shipping a 3-pound package and $40 for shipping a 10-pound package. Now the question is to write an equation that gives the total cost of shipping a package as a function of the weight of the package. I am stuck on this and have no idea how to solve it. Can someone please give me an answer? AND FAST!

OpenStudy (whpalmer4):

price = base price + per pound rate * weight right?

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

okay, let's the base price B, and the per pound rate P then the price to ship a package that weighs x pounds is C = B + Px now, we know two different package weights and their prices. 3 pounds for $22.85 10 pounds for $40 we can make two equations from that data: 22.85 = B + P(3) 40 = B + P(10) can you solve that system of equations for B and P?

OpenStudy (anonymous):

Yes hold on.

OpenStudy (anonymous):

For 22.85 = B + P(3) I got B=P3+22.85. Am I right so far?

OpenStudy (whpalmer4):

if you have 22.85 = B + 3*P and you want B all alone, you can subtract 3*P from both sides: 22.85 - 3*P = B + 3*P - 3*P 22.85 - 3*P = B

OpenStudy (whpalmer4):

do you know about slope intercept form for a line equation?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

do the equations we've written remind you at all of that?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

good! it's exactly the same. in fact, we could solve this problem by making a plot of those two points, drawing a line through them, and seeing where the line crosses the y-axis. that would give us the base price, and the slope of the line is the price per pound.

OpenStudy (whpalmer4):

but let's continue with the approach you were trying. You got B = 22.85-3*P Now substitute that into the other equation: 40 = B + 10*P 40 = 22.85-3*P + 10*P solve that for P

OpenStudy (anonymous):

I got 17.15=P

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

????????

OpenStudy (whpalmer4):

hmm, not quite right. 40=22.85 - 3*P + 10*P 40 = 22.85 + 7P right?

OpenStudy (whpalmer4):

sorry, thought I clicked the Post button, but apparently OS didn't think so :-)

OpenStudy (anonymous):

Oh I made a simple mistake, but I understand now.

OpenStudy (anonymous):

And also a whole bunch of other mistakes.LOL

OpenStudy (whpalmer4):

so what do you get for the value of P? and from that, with the substitution equation, what is the value of B?

OpenStudy (anonymous):

Ok, so I got P=2.45.

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

YAY

OpenStudy (whpalmer4):

so $2.45 per pound. how much is the base price?

OpenStudy (anonymous):

One second...

OpenStudy (whpalmer4):

the cost of shipping a weightless package, if you want to think of it that way :-)

OpenStudy (anonymous):

OK B=15.5?

OpenStudy (whpalmer4):

yep, you nailed it!

OpenStudy (whpalmer4):

should check the work, of course: 3 pound package: C = 15.5 + 3*2.45 = 15.5 + 7.35 = 22.85 10 pound package C = 15.5 + 10*2.45 = 15.5 + 24.50 = 40 looks good!

OpenStudy (anonymous):

K

OpenStudy (whpalmer4):

Here's a graph

OpenStudy (anonymous):

So would that answer the question?

OpenStudy (anonymous):

(the equation that is)

OpenStudy (whpalmer4):

Yes, C = 15.5 + 2.45w if C is the cost and w is the weight

OpenStudy (anonymous):

Thank you, but I have a question about another question. How can you tell if data can be modeled by a linear equation?

OpenStudy (whpalmer4):

well, if you plot the data and it is pretty much a straight line, then a linear equation will model it. Or, even if it isn't a straight line, if the area of interest is small enough, a straight line may model the area of interest adequately.

OpenStudy (anonymous):

I have to write an equation in point-slope form that relates y to x. Can you help me with that?

OpenStudy (whpalmer4):

sure, but you can do it yourself now. just do the same thing we did for the shipping problem!

OpenStudy (anonymous):

Oh ok.

OpenStudy (whpalmer4):

give it a try, I'll answer your questions if you get stuck

OpenStudy (anonymous):

I actually have to get off the computer now, but thank you for your help!

OpenStudy (anonymous):

Your welcome!!

OpenStudy (whpalmer4):

tag me in the future if you need help — just put "@whpalmer4" in a post and I'll get a notification. I'll try to help you when I'm next online.

OpenStudy (anonymous):

Ok I`ll be sure to.